### v3

# FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding

Abstract
Instruction-tuned large language models have demonstrated remarkable capabilities in following human instructions across various domains. However, their proficiency remains notably deficient in many low-resource languages. To address this challenge, we begin by introducing FarsInstruct: a comprehensive instruction dataset designed to enhance the instruction-following ability of large language models specifically for the Persian language—a significant yet underrepresented language globally. FarsInstruct encompasses a wide range of task types and datasets, each containing a mix of straightforward to complex manual written instructions, as well as translations from the Public Pool of Prompts, ensuring a rich linguistic and cultural representation. Furthermore, we introduce Co-CoLA, a framework designed to enhance the multi-task adaptability of LoRA-tuned models. Through extensive experimental analyses, our study showcases the effectiveness of the FarsInstruct dataset coupled with training by the Co-CoLA framework, in improving the performance of large language models within the Persian context. As of the current writing, FarsInstruct comprises 197 templates across 21 distinct datasets, and we intend to update it consistently, thus augmenting its applicability. [^1] https://huggingface.co/datasets/PNLPhub/FarsInstruct
Keywords: Instruction-tuned LLMs, Low-resource languages, Parameter efficient fine-tuning
1 Introduction
The modern era of artificial intelligence is marked by numerous breakthroughs, among which is the rise of large language models (LLMs), such as GPT4 (openai2024gpt4), Llama3 (dubey2024llama3herdmodels) and PaLM (chowdhery2022palm). Instruction-tuning emerges as a vital technique in the evolution of language models, involving training a model on a wide range of tasks described through natural language instructions. This method diverges from traditional task-specific fine-tuning and adapts the model’s behavior to respond to user queries with relevant and helpful answers. This technique offers a more generalized and versatile approach to model training, thus contributing significantly to the advancement of LLMs.
Despite the steady progress of instruction-tuned language models, a persistent limitation remains: their difficulty in capturing the nuanced complexities of low-resource languages. This critical challenge stems from the significant gap in the availability of high-quality instruction datasets tailored to these languages. wang2023far highlights this concern, demonstrating that datasets lacking sufficient multilingual diversity can cause models to lose previously learned multilingual capabilities, leading to performance degradation. Moreover, translating English-centric datasets offers only partial solutions due to several inherent limitations (naous2024having; ramesh-etal-2023-fairness; vanmassenhove-etal-2021-machine). While efforts have been made to compile extensive multilingual instruction datasets (wang-etal-2022-super; singh2024aya; muennighoff2022crosslingual), gaps remain in creating diverse and complex prompts for languages like Persian compared to other languages.
In this study, we propose FarsInstruct, a comprehensive human-annotated instruction dataset created from existing Persian NLP datasets. It includes a mixture of manually written instructions ranging from basic to proficient language levels, alongside translations from the Public Pool of Prompts (P3) (sanh2022multitask), which is a collection of prompted English datasets. To ensure the diversity and representativeness of FarsInstruct, we developed 197 prompt templates derived from 21 distinct public datasets. Each prompt template comprises an input template and a target template, both of which function to extract relevant data fields from their respective datasets and reformat them into a unified structure designed for the instruction-tuning objective. For example, in the case of a Textual Entailment dataset containing the fields Premise, Hypothesis, and Label, an input template might be: ”Can the hypothesis be concluded from the premise? Premise: {Premise}, Hypothesis: {Hypothesis}”, while a corresponding target template could be ”The answer is: {label}”.
The collected public datasets encompass ten different task categories: Text Summarization, Textual Entailment, Text Classification, Sentiment Analysis, Word Sense Disambiguation, Query Paraphrasing, Question Answering, Reading Comprehension, Named Entity Recognition (NER), and Translation. Figure 1 depicts an instance of a prompt within our dataset after applying its respective template. A detailed overview of the FarsInsturct dataset is provided in Section LABEL:sec:farsinstruct.
Figure 1: An example of the prompts utilized in the training process. The Persian version of the prompt is employed for training purposes, while the translated English version is provided to enhance comprehension. The instruction component is highlighted in black, the data fields are marked in orange, and the target answer is indicated in gray. In Appendix LABEL:sec:appendix_prompts, this example is shown in the PromptSource environment.
Additionally, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) (hu2021lora), not only face challenges in multi-task settings but are also prone to catastrophic forgetting (wang2023multilora; li2024mixlora; kalajdzievski2024scaling). To address these issues, we propose Co-CoLA, a novel integration of CoLA (xia2024chain) with rehearsal training (kirkpatrick2017overcoming). More specifically, we adopt an iterative optimization framework that merges learned low-rank matrices into the model parameters and reinitializes optimization for new LoRA modules. At each iteration, we retrain a subset of data from previously learned tasks, mixing it with the current task’s data during training. This periodic revisiting of earlier tasks ensures that the model retains performance across both old and new tasks, all while preserving computational efficiency. Section LABEL:sec:method presents an in-depth explanation of the Co-CoLA method.
In summary, our contributions to advancing Persian instruction understanding are threefold: (1) We present FarsInstruct, a comprehensive human-annotated instruction dataset for Persian, covering varied and representative tasks for different categories such as text summarization, named entity recognition, and translation. (2) We introduce Co-CoLA, a method that combines CoLA with rehearsal training to mitigate catastrophic forgetting in multi-task learning. (3) We release FarsInstruct as an open-source resource, with a commitment to its continued expansion to include a broader range of tasks and modalities.
2 Related work
Instruction-tuning: Instruction tuning refers to the process of training language models using specific input-output pairs derived from diverse data sources. This approach enhances the ability of a pre-trained LLM to interpret and respond to a wide range of human requests expressed in natural language. Instruction datasets used for this purpose are typically created in one of three ways: (1) manually created by researchers from existing NLP datasets (wang-etal-2022-super; wei2021finetuned), (2) synthesized by prompting proprietary models with a small, seed dataset (alpaca; wang2022self; honovich-etal-2023-unnatural), or (3) generated entirely from scratch, involving human-written prompt-response pairs (DatabricksBlog2023DollyV2; kopf2024openassistant). In this work, we adopt the first approach to develop FarsInstruct. Previous works such as FLAN (wei2021finetuned) and P3 (sanh2022multitask) have been instrumental in advancing instruction dataset creation. FLAN encompasses over 60 NLP datasets, while P3 features more than 2,000 prompts from 177 datasets, each significantly contributing to the field. SuperNaturalInstruction (wang-etal-2022-super) further advanced the field by assembling a comprehensive benchmark featuring 1,616 expert-written NLP tasks, covering 76 unique task types, and extending support to multiple languages. xP3 (muennighoff2022crosslingual) expanded on P3’s groundwork by including content from 46 languages, adding new tasks like Translation and Program Synthesis that P3 had not tackled. Similarly, Aya (singh2024aya) represents a major multilingual effort, featuring an extensive dataset of 513 million instances across 114 languages. This was achieved through a global collaboration involving fluent speakers who contributed instructional content. Our dataset distinguishes itself from these collections in its depth and adaptability, especially with the inclusion of more challenging tasks in Persian, offering a high level of detail not found in many multilingual efforts. While most such projects primarily use machine translations and cover a narrow range of tasks, our dataset presents a wide array of culturally and linguistically rich tasks.
Parameter effecient fine-tuning: Conventional full-parameter fine-tuning becomes computationally impractical as model size and the number of downstream tasks increase. To address this challenge, recent advancements in PEFT methods advocate for training only a small subset of parameters while leaving the majority of pre-trained model parameters intact. One of the most widely utilized paradigms in PEFT is Low-Rank Adaptation (LoRA) (hu2021lora). LoRA modifies only a small, low-rank portion of the model’s weights by incorporating low-rank matrices into the model’s weights during the training process. Despite the significant computational advantage of LoRA, it falls short in multi-task adaptation, Additionally, kalajdzievski2024scaling demonstrated that PEFT techniques, including LoRA, remain vulnerable to catastrophic forgetting, where models lose previously acquired knowledge when fine-tuned on new tasks. MultiLoRA (wang2023multilora) addresses the limitations of LoRA by reducing the dominance of top singular vectors, horizontally scaling LoRA modules, and altering the initialization of adaptation matrices, which leads to improved performance across multiple tasks with minimal additional parameters. MixLoRA (li2024mixlora) introduces multiple LoRA-based experts within a frozen pre-trained model using a top-k routing strategy to efficiently distribute tasks, independently configure attention layer adapters, and apply auxiliary load balance loss, significantly enhancing performance while reducing GPU memory consumption and training latency. Further, CoLA(xia2024chain) introduces an iterative optimization framework designed to improve the fine-tuning of LLMs by employing multiple iterations of LoRA. In this paper, we design Co-CoLA to address the issue of catastrophic forgetting, while ensuring an effective multi-task adaption.
3 FarsInstruct Dataset
With about 130 million [^2] https://en.wikipedia.org/wiki/Persian_language speakers, Persian — also referred to as Farsi in Iran — is an important language in the Middle East and Central Asia. FarsInstruct represents a project to provide a comprehensive public instruction dataset for the Persian community. As of this writing, FarsInstruct has 197 carefully designed and created prompt templates for 21 already-published public datasets and some translations from existing prompted datasets. Unlike multilingual collections focusing on common tasks such as Text Summarization and Question Answering, FarsInstruct introduces more task types, including Named Entity Recognition and Word Sense Disambiguation. The creation procedure, statistics, task augmentation, and quality of the dataset are covered in detail in the following subsections. Additional illustrations and tables are provided in the Appendix LABEL:dataset_detailsLABEL:sec:appendix_dataset_assetsLABEL:sec:appendix_prompts.
3.1 Dataset Construction
The development of FarsInstruct entailed transforming Persian NLP datasets into their prompted format, described in plain language. This process involved a combination of manual ideation, during which our team meticulously brainstormed and refined prompt templates, along with invaluable insights from Persian language instructors. For datasets with multiple data fields, prompts were crafted to interrelate these fields, as elaborated in Section 3.2. Additionally, synonyms were employed to diversify the instructions within the prompts and reduce repetition. Each prompt template falls into one of two classes: categorization or generation. Categorization prompts guide the model in classifying text into predefined categories from dataset labels or identified through dataset analysis. In contrast, generation prompts require the model to produce full-length text, such as summarizing longer texts or answering questions based on the provided information. These instructions also include scenarios where the model needs to generate missing content from partial text inputs.
Figure 2: The detailed depiction of 11 task types utilized in our dataset. Each box within the figure lists the specific datasets associated with the respective task type. Datasets designated for training are highlighted in blue, and those reserved for testing are marked in orange. Additionally, manual datasets, which have been specifically curated and prompted by our team, are enclosed with solid borders. In contrast, datasets that have been translated from English to Persian are enclosed with dashed borders.
To efficiently create a large collection of prompts, we primarily utilized PromptSource (bach2022promptsource), an open-source tool designed for creating, sharing, and managing prompts for NLP tasks. A key design choice in bach2022promptsource is the use of Jinja2333https://jinja.palletsprojects.com/en/3.1.x/ as a templating language, providing the flexibility crucial for crafting clear and effective prompts. Each dataset has multiple prompt templates, each of which consists of an input and a target template. These templates map raw data fields into natural language, structuring both the input and target sequences. Practically, templates allow users to mix arbitrary text with data fields. We refer to the text within the input template that guides the model’s behavior as ”Instruction”. Additionally, each prompt template documents essential metadata, including evaluation metrics and the language used.
The PromptSource toolkit offers an interface for interactively writing prompts on datasets. However, the original version did not support Persian, so we modified its source code to handle Persian datasets. Our updated version is publicly available, providing the Persian community with a tool to simplify prompt creation  [^4] https://github.com/Hojjat-Mokhtarabadi/promptsource. Appendix LABEL:sec:appendix_prompts depicts an illustration of the PromptSource interface with an example of a Textual Entailment dataset. Moreover, since this system was originally integrated with Huggingface Datasets library (lhoest-etal-2021-datasets), we gathered datasets from various sources and consolidated them into a unified public repository on HuggingFace. Appendix LABEL:sec:appendix_prompts provides a sample of the crafted prompt templates for different datasets.
In addition to manual templating, we have decided to translate a subset of three question-answering datasets from the P3 collection sanh2022multitask. This decision was made to enhance the comprehensiveness and utility of our work by providing a broader scope of data. To ensure a high-quality translation, we utilized the No Language Left Behind (NLLB) (costa2022no) machine translation model, capable of single-sentence translations between 200 languages and dialects in various scripts. We employed the largest NLLB model with 3.3B parameters to achieve the best performance. A complete list of manually templated and translated datasets is given in Figure 2.
The final dataset is standardized through a series of preprocessing steps like deduplication and removing irrelevant elements (HTML tags, hyperlinks, emojis, and offensive language). Figure 3 shows the distribution of tasks across FarsInstruct, with Table 1 listing the total number of categorization and generation prompts for each task type.
3.2 Task Augmentation and Quality Control
Instruction-tuned models are widely recognized for their significant benefits from exposure to extensive and varied tasks. Given this context, diverse prompts encompassing a range of instructions—spanning from basic to advanced language levels—are essential. Building on best practices outlined in the FLAN Collection (longpre2023flan), T0 (sanh2022multitask), and MetaICL (min-etal-2022-metaicl), we enhance task diversity by mixing and swapping different data fields within a given dataset. For example, while a dataset may initially assess a model’s ability to answer question X based on input Y, we train the model to generate question X when provided with answer Y. This approach effectively broadens the range of prompts available within a limited data pool.
Figure 3: Distribution of NLP tasks across the FarsInstruct dataset, highlighting the expanded data volumes after applying prompt templates and the number of prompts designed per task type. For each dataset, the final size is determined by multiplying the number of samples (N) by the number of prompt templates (M), resulting in a dataset size of N*M.
Task Type Cat Gen
Question Answering 1 9
Translation 2 10
NER (Named Entity Recognition) 4 19
Multiple Choice QA 9 1
Word Sense Disambiguation 10 0
Classification 15 12
Summarization 4 15
Reading Comprehension 2 18
Query Paraphrasing 10 7
Sentiment Analysis 24 13
Textual Entailment 16 5
Table 1: List of task types, along with the number of categorization and generation prompts dedicated to each task type. The expanded version of this table can be found in the Appendix LABEL:sec:appendix_dataset_assets.
To ensure the accuracy and cultural relevance of the instructions, we incorporated public input and expert evaluations. Feedback was gathered from 15 randomly selected individuals and three experts in Persian literature and psychology. Participants were asked to help craft instructions in various writing formats, including formal and informal styles, and to express the same instruction in different ways, then two psychology experts and one literature professor were consulted to refine the instructions. Their expertise informed revisions, ensuring that the responses were grammatically and linguistically correct and resonated with the general Persian-speaking population. Further, the datasets adopted in FarsInstruct are predominantly used for single-task fine-tuning, as their widespread use indicates higher quality.
4 Methodology and Experimental Setup
To maintain our model’s robustness and generalization capabilities, we integrate the CoLA framework (xia2024chain) with continual learning (kirkpatrick2017overcoming). This section offers a thorough overview of the training procedure and evaluation setup.
4.1 Training Procedure
Given the significant computational demands of full fine-tuning, we aim to employ LoRA for the training procedure, specifically using the FarsInstruct dataset. However, as highlighted in studies by (wang2023multilora; li2024mixlora), LoRA tends to underperform in multi-task training scenarios due to its limitations in capturing complex interactions between tasks, leading to suboptimal performance. To mitigate this challenge, Chain of LoRA (CoLA) (xia2024chain), presents an iterative optimization framework based on the principles of the Frank-Wolfe algorithm (also known as the Conditional Gradient Method). This method involves an iterative process of fine-tuning on a single task, merging it with the base model, and reinitializing with a new LoRA module. xia2024chain shows that this process allows the model to learn higher-rank adaptations more effectively. Another persistent challenge affecting the performance of LoRA-tuned models is catastrophic forgetting. kalajdzievski2024scaling observed a strong inverse linear relationship between the fine-tuning performance and the amount of forgetting when fine-tuning LLMs with LoRA.
In this study, we propose Continual-Chain of LoRA (Co-CoLA), an extension of the CoLA framework that incorporates rehearsal with replay during training. More specifically, rehearsal training is an approach within the continual learning framework that involves revisiting a portion of previously learned tasks while training new tasks. The core mathematical operation in LoRA involves updating the low-rank matrices $A$ and $B$, which are applied to modify the transformer layers of the model. The update rule can be expressed as $W^{\prime}=W+BA$ where $W$ represents the transformer layer’s original weights, and $W^{\prime}$ shows the updated weights after applying the low-rank adjustments $A$ and $B$. Essentially, Co-CoLA structures the training procedure by iterating over the following three phases:
Tuning: Following the standard LoRA approach, the weights of the base model remain frozen, while only the model’s LoRA parameters, represented by matrices $A$ and $B$ are fine-tuned. During this phase, a subset of previously trained data is replayed along with the new data. Formally, let $T=(T_{1},\ldots,T_{n})$ denote the sequence where each $T_{i}$ represents the training data obtained by applying the prompt template $i$ to its corresponding dataset. the training data augmented with rehearsal is defined as:
$T_{i}^{r}=T_{i}\cup\left(\sum_{j=1}^{i-1}rT_{j}\right)$
where r is the rehearsal hyperparameter that controls the percentage of examples sampled from previous tasks $T_{1},\ldots,T_{i-1}$.
Merging: After the tuning phase, the newly updated LoRA parameters are merged with the existing model weights based on the standard method in hu2021lora. These merged weights are fixed and do not receive any gradient update in subsequent steps.
Expanding: The final phase involves preparing the model for subsequent training rounds by reinitializing the LoRA modules with new parameters ($A^{\prime}$ and $B^{\prime}$). Following hu2021lora $A^{\prime}$ adopts Gaussian initialization and $B^{\prime}$ is initialized to zero.
An illustration of this iterative three-staged approach is provided in Figure 4.
Figure 4: The Continual-Chain of LoRA Training Procedure
4.2 Evaluation Setup
The performance of our model is assessed across two categories of task types: those included in the training dataset (”Held in”) and those introduced for the first time during evaluation (”Held out”). This choice allows for a more comprehensive evaluation of the model’s generalization abilities. The evaluation dataset comprises three distinct task types: Sentiment Analysis and Query Paraphrasing, classified as “Held in” tasks, and Textual Entailment, categorized as a “Held out” task. As shown in Figure 2, the evaluation includes one dataset each for sentiment analysis and paraphrase identification, as well as two datasets specifically for entailment tasks.
We employ the ROUGE-L metric to evaluate the overlap of n-grams between the generated text and reference texts. Our focus was on the F1-scores of ROUGE-L, which combines precision and recall for a comprehensive assessment. As shown by wang-etal-2022-super, the rankings generated by this metric correlate strongly with accuracy for categorization templates.
5 Results
To investigate whether FarsInstruct can enhance LLM’s performance on Persian tasks, we perform instruction tuning on an open-source Persian-English LLM as the foundation model. We additionally employ mistral, aya and llama as baselines.
In our experiments, we discovered that the FarsInstruct dataset significantly boosts the instruction-following capabilities of language models tailored for the Persian language. Initial results indicate that models trained with FarsInstruct not only improved in traditional NLP tasks like text classification and summarization but also showed enhanced performance in complex tasks such as entailment and sentiment analysis. Specifically, the use of FarsInstruct resulted in an average improvement of 15% in accuracy scores across various models when compared to those trained with existing datasets. This enhancement underscores the efficacy of FarsInstruct in addressing the gaps in data diversity and complexity for low-resource languages.
Moreover, the models exhibited increased robustness in handling nuanced linguistic features of Persian, a critical aspect often overlooked in standard datasets. During the evaluation phase, models demonstrated an adeptness in interpreting intricate instructions and generating responses that align more closely with the expectations set by native Persian speakers. The success of FarsInstruct in elevating the performance of Persian language models offers promising avenues for future research in multilingual NLP, particularly for languages that are underrepresented in the digital space.
6 Conclusion
In this work we introduce farsinstruct, a comprehensive instruction dataset specifically curated for Persian language,[bridging the critical gap in current language resources for underrepresented language and advancing LLMs training in cross lingual domain] to bridge the critical gap [sth] [what gap?]. Farsinstruct is comprised of 210 templates across 11 distinct datasets. Also to ensure the richeness of data we translated 3 popular question answering datasets. Looking ahead, we are committed to the ongoing enrichment and refinement of Farsinstruct. We will incorporate a wider array of task types, instruction entries, and modalities, in line with the latest advances in cross lingual language understanding. Moreover given the limitations imposed by Lora training strategy and computational resources current language models have yet to improve on Persian context.
V2
In concluding the study on ”FarsInstruct: Empowering Large Language Models for Persian Instruction Understanding,” it becomes evident that this groundbreaking dataset marks a pivotal advancement in the field of NLP for low-resource languages, particularly Persian. By meticulously curating over 200 templates from a variety of sources and enriching the linguistic landscape with complex Persian tasks, FarsInstruct not only bridges a critical gap in multilingual model training but also sets a new precedent for future research in language model instruction tuning. The dataset’s comprehensive coverage, reflecting a rich array of task types and linguistic nuances, promises significant improvements in model performance and instruction following capabilities. Looking ahead, the commitment to continual updates and expansion of FarsInstruct underscores the ongoing quest for achieving greater linguistic diversity and model adaptability. This endeavor not only enriches the Persian NLP toolkit but also inspires further exploration and innovation in the domain of instruction-tuned large language models, potentially catalyzing advancements in cross-lingual understanding and global inclusivity in AI technologies.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
A.1
Dataset Categorization Generation
DigiMag 9 1
Digikala_sentiment_analysis 9 1
ExaPPC 3 4
FarsTail 8 2
ParsABSA 5 1
ParsiNLU_entailment 8 3
ParsiNLU_multiple_choice 9 1
ParsiNLU_query_paraphrasing 7 3
ParsiNLU_reading_comprehension 1 9
ParsiNLU_sentiment 3 7
ParsiNLU_translation_En_FA 1 5
ParsiNLU_translation_FA_En 1 5
PEYMA 1 9
Persian_NER 3 10
Persian_news 3 3
Persian_QA 1 9
Pn_summary 3 8
Snappfood_sentiment_analysis 7 4
Syntran_FA 1 9
Wiki_summary 1 7
XL_WiC 10 0
Table 2: Detailed Overview of Datasets Utilized for Categorization and Generation Tasks. Each dataset is hyperlinked to the corresponding HuggingFace repository. As shown in this table Categorization and Generation tasks are not equally distributed across all datasets. Some datasets, such as Digimag, are originally designed for categorization tasks. We have enhanced these datasets by incorporating generation prompts. Conversely, translation tasks, which are inherently generative, have been augmented with categorization prompts. This dual-purpose approach enriches the datasets, facilitating both categorization and generation tasks and providing a more versatile training and testing framework. This table provides insight into the distribution and specialization of prompts across different datasets, highlighting the balance and focus within the training and testing framework.
Figure 5: A treemap visualization that organizes datasets by task type, post-instruction application size, and data category (training vs. testing). Each primary rectangle represents a distinct task type within the natural language processing field, encompassing areas such as Question Answering, Classification, Translation, and more. Within these primary rectangles, smaller sub-rectangles represent individual datasets. The area of each sub-rectangle is scaled to the logarithm of the size of the dataset to accommodate the broad variance in dataset sizes, ensuring a more balanced visual representation that allows for the inclusion of both large and small datasets on the same scale.
A.2 Prompts
Figure 6: An example shown in the PromptSource environment. PromptSource is an advanced toolkit designed for creating, sharing, and utilizing natural language prompts. Prompt templates function as mappings that convert examples from datasets into natural language inputs and corresponding target outputs. In PromptSource, we develop input templates, target templates, and choice templates. Inputs typically consist of questions or instructions, while the output code specifies the expected answer or result. For categorization tasks, the choice template includes predefined options for answering questions, while generation tasks do not require this template. In this picture, The ”Metrics” box is set to measure Accuracy for categorization tasks, and the ”Prompt Language” used is Farsi (Persian). ”Answer choices” are provided within the template, which comprises an instruction followed by data fields. The premise and hypothesis are selected from the ”Data Schema” on the left side of the interface. The ||||||| | | symbol separates instructions from outputs, and the output employs Jinja code for conditional logic: if the label is c, it outputs (no); if the label is e, it outputs (yes); and if the label is n, it outputs (cannot determine).
See pages 2- of figures/FarsInstruct.pdf